### v1

# Non-local Metric-Affine Gravity

Abstract
Non-local gravity can potentially solve several problems of gravitational field both at Ultra-Violet and Infra-Red scales. However, such an approach has been formulated mainly in metric formalism. In this paper, we discuss non-local theories of gravity in the metric-affine framework. In particular, we study the dynamics of metric-affine analogue of some well-studied non-local theories, by treating the metric and the connection as independent fields. The approach gives the opportunity to deal with non-local gravity under a more general standard. Furthermore, we introduce some novel non-local metric-affine theories with no Riemannian analogue and investigate their dynamics. Finally we discuss some cosmological applications of our development.
I Introduction
It is widely believed that quantum gravitational effects manifest through non-locality. This feature is invoked in order to reconcile General Relativity (GR) with Quantum Field Theory (QFT) prescriptions. These quantum corrections, that alter the classical gravitational action, encode non-locality and dramatically change the associated dynamics producing new degrees of freedom. Non-locality is typically expressed through the inclusion of inverse powers of the box operator $\Box^{-1}$ sandwiched between geometric quantities (like the Ricci scalar) that supplement the gravitational action and consequently the associated field equations Mashhoon:2017qyw ; Deser:2007jk ; Capozziello:2008gu ; Nojiri:2010pw . Another kind of non-locality is realized by adding an infinite series of the box operator in the form of entire functions like e.g. $RF(\Box)R$ where $R$ is the Ricci scalar and $F(\Box)$ a holomorphic function of the box operator. This is known as Infinite Derivative Gravity (IDG) efimov1967non . Besides fundamental physics issues, the key idea to adopt non-local gravity models is to obtain characteristic lengths and masses capable of addressing Ultra-Violet and Infra-Red issues in astrophysics and cosmology Blome:2010xn ; Roshan:2022zov .
Cosmological applications of non-local gravity have been studied in Refs. Deser:2007jk ; Koivisto:2008xfa ; Calcagni:2018pro ; Capozziello:2024qol ; Capozziello:2021krv ; Nojiri:2019dio ; Biswas:2010zk ; Kolar:2021qox ; Koshelev:2016xqb ; Calcagni:2013vra ; Dirian:2014bma and some problems like the Weinberg no-go theorem for cosmological constant can be addressed in this framework Capozziello:2025bsm . The teleparallel formulation of non-local gravity along with its cosmological applications, has been considered in Bahamonde:2017bps . Also black hole solutions can be improved with the presence of non-local terms Bambi:2016uda ; Capozziello:2025iwn as well as gravitational waves where new polarizations emerge Capozziello:2024bjb ; Foffa:2013sma .
Here we are going to formulate non-local gravity in the metric-affine geometry setup in order to consider the most general framework where these theories can be formulated. This approach proves to be useful also in view of extensions and modifications of General Relativity where the geometric invariants can be other than the standard curvature Bahamonde:2017bps ; BeltranJimenez:2019esp ; Capozziello:2022zzh ; Capozziello:2025hyw ; Golovnev:2023yla .
The paper is organized as follows. In Sec.IIII, we introduce the notation and conventions we are going to use. Then, in Sec. $III$, we start with the metric-affine analogue of the theory introduced in Ref. Gottlober:1989ww , which was studied in the Riemannian setting. Immediately the differences with the metric formalism emerge. By solving the connection field equations of the theory we cast it in an equivalent Scalar-Tensor metric theory with higher order derivative terms for the scalar field. Consequently we generalize it in Sec. $IV$ and study the metric-affine version of the general class of theories with the Lagrangians $\mathcal{L}=RF(\Box)R$ and $\mathcal{L}=RF(\Box^{-1})R$. Deriving the field equations, and solving the connection equations, we again establish equivalence with a generalized higher-derivative Scalar-Tensor theory over a Riemannian background. Finally, we introduce a new non-local Metric-Affine Gravity theory which has no Riemannian analogue and consists of the newly introduced terms $X\Box X$, where $X$ stands for torsion and non-metricity, into the action, along with the usual Einstein-Hilbert term. Obtaining the field equations we show how torsion and non-metricity are dynamical fields in this setting even in vacuum. Some applications to cosmology are discussed in Sec $VI$. In Sec. $VII$, we then wrap up our results and point to future directions.
II The Metric-Affine formalism
Let us consider a $n$-dim manifold equipped with a metric $g$ and a linear connection $\nabla$, whose components, in local coordinates read $g_{\mu\nu}$ and $\Gamma^{\lambda}{}_{\mu\nu}$ respectively. With these at hand, we can accordingly define curvature, torsion and non-metricity, respectively, as
$R^{\mu}_{\;\;\;\nu\alpha\beta}$ $:=2\partial_{[\alpha}\Gamma^{\mu}_{\;\;\;|\nu|\beta]}+2\Gamma^{\mu}_{\;\;\;\rho[\alpha}\Gamma^{\rho}_{\;\;\;|\nu|\beta]}\;\;,$ (1a)
$S_{\mu\nu}^{\;\;\;\lambda}$ $:=\Gamma^{\lambda}_{\;\;\;[\mu\nu]}\;\;,\;\;$ (1b)
$Q_{\alpha\mu\nu}$ $:=-\nabla_{\alpha}g_{\mu\nu}$ (1c)
The Ricci tensor for the general connection is given by the contraction
$R_{\mu\nu}:=R^{\alpha}_{\;\;\;\mu\alpha\nu}$ (2)
It should be noted that this is not symmetric in general. Subsequent contraction with the metric defines the generalized scalar curvature
$R:=g^{\mu\nu}R_{\mu\nu}$ (3)
Now, from torsion and non-metricity, one can construct the three vectors
$S_{\mu}:=S_{\mu\nu}{}{}^{\nu}\;\;,\;\;Q_{\mu}:=Q_{\mu\alpha\beta}g^{\alpha\beta}\;\;,\;\;q_{\mu}:=Q_{\alpha\beta\mu}g^{\alpha\beta}$ (4)
The difference between the general connection $\Gamma^{\lambda}{}_{\mu\nu}$ and the Levi-Civita one, $\tilde{\Gamma}^{\lambda}{}_{\mu\nu}$, defines a tensor that quantifies how much the geometry deviates from being Riemannian. This is the so-called distortion tensor schouten1954ricci ; hehl1995metric
$N^{\lambda}{}_{\mu\nu}:=\Gamma^{\lambda}{}_{\mu\nu}-\tilde{\Gamma}^{\lambda}{}_{\mu\nu}$ (5)
Then, using the above definitions of torsion and non-metricity, it is a simple matter to show that the general affine-connection may be decomposed according to
$\Gamma^{\lambda}_{\;\;\;\;\mu\nu}=\tilde{\Gamma}^{\lambda}_{\;\;\;\mu\nu}+\frac{1}{2}g^{\alpha\lambda}(Q_{\mu\nu\alpha}+Q_{\nu\alpha\mu}-Q_{\alpha\mu\nu})-g^{\alpha\lambda}(S_{\alpha\mu\nu}+S_{\alpha\nu\mu}-S_{\mu\nu\alpha})$ (6)
This is called the post-Riemannian expansion of the connection and it is particularly useful for splitting the various quantities into their Riemannian parts [^1] Namely the ones computed with respect to the Levi-Civita connection. plus the non-Riemannian contributions of torsion and non-metricity. Torsion and non-metricity can then be recovered from the distortion from the relations (see e.g. Hehl:1994ue ),
$Q_{\mu\alpha\beta}=2N_{(\alpha\beta)\mu}\;\;,\;\;S_{\mu\nu\alpha}=N_{\alpha[\mu\nu]}.$ (7)
Our definition of the box operator is the usual one that involves only the metric, i.e.
$\Box f:=\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}\partial^{\mu}f)$ (8)
on a test function $f$. In a covariant manner, it can also be written as
$\Box=g^{\mu\nu}\widetilde{\nabla}_{\mu}\widetilde{\nabla}_{\nu}=\widetilde{\nabla}^{\mu}\widetilde{\nabla}_{\mu}\,.$ (9)
In this form, it can further act upon tensor fields of arbitrary rank, $\Box T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}=\widetilde{\nabla}^{\mu}\widetilde{\nabla}_{\mu}T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}$. Note that in Briscese:2015urz the definition
$\Box^{(\Gamma)}=g^{\mu\nu}\nabla_{\mu}\nabla_{\nu},$ (10)
where $\nabla$ is the full covariant derivative associated with the full connection $\Gamma^{\lambda}{}_{\mu\nu}$, was used in a Palatini formulation of non-local gravity. Although this could be a viable option, there are several reasons to avoid such a definition in the current context. First and foremost, (10) does not reduce to (8) when acted upon functions. Secondly, given the fact that the metric is not covariantly constant with respect to the full connection, the above definition is not the only one that reduces to $\partial^{\mu}\partial_{\mu}$ in the flat space limit. Indeed, operators like $\nabla_{\mu}(g^{\mu\nu}\nabla_{\nu})$, $\frac{1}{\sqrt{-g}}\nabla_{\mu}(\sqrt{-g}g^{\mu\nu}\nabla_{\nu})$ or $\nabla_{\mu}(\nabla_{\nu}g^{\mu\nu})$ also reduce to $\partial^{\mu}\partial_{\nu}$ in the flat space limit. One would then need to take all possible linear combinations of such operators (with the coefficients added up to one) in order not to discriminate among them, see for instance Shimada:2018lnm . However, this approach, not only does it complicate the analysis but also introduces new arbitrary parameters. Thirdly, for operators like (10), terms like $R_{\mu\nu}\Box^{(\Gamma)}R^{\mu\nu}$ and $R^{\mu\nu}\Box^{(\Gamma)}R_{\mu\nu}$ do not coincide since the metric cannot freely move through the full covariant derivative $\nabla_{\mu}$. This then complicates the analysis even further since one would be forced to include both of them in order to maintain full generality. Finally, the basic property of the inner product $(\phi\Box\psi)=(\Box\phi,\psi)$ is lost if one uses (10) instead of (8). It is therefore only natural to stick to (8) for the definition of the box operator in our current context.
III A simple Warm-up Theory
Let us start with the model in Gottlober:1989ww , but now in the metric-affine formalism, where the connection and the metric are independent fields. We have
$S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(\alpha R+\beta R^{2}+\gamma R\Box R\Big)$ (11)
where $\alpha$ is dimensionless [^2] Of course, the General Relativity limit is recovered for $\alpha=1$. and $\beta$ and $\gamma$ are parameters with mass dimension 2-2 and 4-4 respectively. Also, we define the box operator associated to the metric $g$, in the usual way, (see eq. (8)) as we explained in the previous section. Let us also note that, in this action, the scalar curvature $R$ is now the generic one that has also torsion and non-metricity. The action (11) is, up to surface terms, equivalent to
$S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(\alpha R+\beta R^{2}-\gamma g^{\mu\nu}\partial_{\mu}R\partial_{\nu}R\Big)\,.$ (12)
We shall use the latter form since it is more convenient to perform the variations. Varying with respect to the metric and the independent affine-connection, we obtain the field equations of the non-local theory
$\frac{1}{2}\Big(\alpha R+\beta R^{2}-\gamma(\partial R)^{2}\Big)g_{\mu\nu}+\Big(\alpha+2\beta R+2\gamma\Box R\Big)R_{(\mu\nu)}-\gamma\partial_{\mu}R\partial_{\nu}R=0$ (13)
and
$\phi P_{\lambda}{}^{\mu\nu}+\delta_{\lambda}^{\nu}\partial^{\mu}\phi-g^{\mu\nu}\partial_{\lambda}\phi=0$ (14)
respectively, where
$P_{\lambda}^{\;\;\;\mu\nu}=-\frac{\nabla_{\lambda}(\sqrt{-g}g^{\mu\nu})}{\sqrt{-g}}+\frac{\nabla_{\sigma}(\sqrt{-g}g^{\mu\sigma})\delta^{\nu}_{\lambda}}{\sqrt{-g}}+2(S_{\lambda}g^{\mu\nu}-S^{\mu}\delta_{\lambda}^{\nu}+g^{\mu\sigma}S_{\sigma\lambda}^{\;\;\;\;\nu})$
$=\delta^{\nu}_{\lambda}\left(q^{\mu}-\frac{1}{2}Q^{\mu}-2S^{\mu}\right)+g^{\mu\nu}\left(\frac{1}{2}Q_{\lambda}+2S_{\lambda}\right)-(Q_{\lambda}^{\;\;\;\mu\nu}+2S_{\lambda}^{\;\;\;\;\mu\nu})$ (15)
is the so-called Palatini tensor and we have abbreviated
$(\partial R)^{2}=g^{\mu\nu}\partial_{\mu}R\partial_{\nu}R\;\;,\;\;\phi=\alpha+2\beta R+2\gamma\Box R$ (16)
Using the method developed in Iosifidis:2018jwu , from the above connection field equations, one finds the distortion
$N_{\alpha\mu\nu}=\frac{2}{(n-2)\phi}g_{\nu[\alpha}\partial_{\mu]}\phi+\frac{1}{2}g_{\alpha\mu}q_{\nu}$ (17)
Now notice that our action (11) is invariant under projective transformations of the connection
$\Gamma^{\lambda}{}_{\mu\nu}\mapsto\Gamma^{\lambda}{}_{\mu\nu}+\delta^{\lambda}_{\mu}\xi_{\nu}$ (18)
since it is built solely from $R$ which is itself projectively invariant. This means that there is an undetermined vectorial degree of freedom which can be fixed by a gauge choice iosifidis2019exactly . Two special gauge choices are of interest here. The first one is to fix
$\xi_{\nu}=-\frac{1}{2}q_{\nu}$ (19)
in which case, as can be easily seen from (7), the full non-metricity vanishes
$Q_{\alpha\mu\nu}=0$ (20)
and torsion is given by
$S_{\mu\nu\alpha}=\frac{1}{(n-2)\phi}g_{\alpha[\nu}\partial_{\mu]}\phi$ (21)
We call this the zero non-metricity gauge. On the other hand, if we fix the gauge as
$\xi_{\nu}=-\frac{1}{2}q_{\nu}+\frac{1}{(n-2)}\frac{\partial_{\nu}\phi}{\phi}$ (22)
then the full torsion is found to be zero, i.e.
$S_{\mu\nu\alpha}=0$ (23)
and the non-metricity takes the simple Weyl form
$Q_{\alpha\mu\nu}=\frac{1}{(n-2)\phi}(\partial_{\alpha}\phi)g_{\mu\nu}$ (24)
This is the vanishing torsion gauge. We see therefore a torsion/non-metricity duality similar to the one that appears in $f(R)$ theories Iosifidis:2018zjj . Let us now investigate the metric field equations (13). Taking the trace of the latter, it follows that
$\left(\frac{n}{2}-1\right)\Big(\gamma(\partial R)^{2}-\alpha R\Big)+2\beta\left(1-\frac{n}{4}\right)R^{2}+2\gamma R\Box R=0$ (25)
The last describes the propagation of a scalar mode. From now on let us set $\alpha=1$, $\beta=0$ and $n=4$ and also denote the scalar mode as $R\equiv\psi$. Substituting the trace eq. (25) back in (13), the latter simplifies to
$R_{(\mu\nu)}-\frac{R}{2}g_{\mu\nu}=\frac{1}{\phi}\left(\gamma\partial_{\mu}\psi\partial_{\nu}\psi-\frac{\psi}{2}g_{\mu\nu}\right)$ (26)
Performing a post-Riemannian expansion on the quantities of the left-hand side of the above, the latter takes the form (see appendix for details)
$\phi\left(\tilde{R}_{\mu\nu}-\frac{\tilde{R}}{2}g_{\mu\nu}\right)=\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\phi-\frac{\partial_{\mu}\phi\partial_{\nu}\phi}{\phi}-\Big[\Box\phi-\frac{3}{4}\frac{(\partial\phi)^{2}}{\phi}\Big]g_{\mu\nu}+\gamma\partial_{\mu}\psi\partial_{\nu}\psi-\frac{\psi}{2}g_{\mu\nu}$ (27)
with $\phi=1+2\gamma\Box\psi$. We see therefore that the initial theory is on-shell equivalent to a Scalar-Tensor theory with higher order derivatives for the scalar field. It is important to stress that, in constrast to the Riemannian case where the resulting metric field equations contain sixth order derivatives of the scalar field, in the metric-affine approach developed here, we have up to fourth-order derivatives of the scalar field. In order words, with respect to the metric formalism developed in Gottlober:1989ww , the metric-affine approach reduces the dynamics. See also Amendola:1993bg for a cosmological application.
IV Non-local Metric-Affine Theories
IV.1 Nonlocal Ultraviolet $\mathcal{L}=R+RF(\Box)R$ Theories
We shall now generalize the results of the previous section and consider the Non-local Metric-Affine Gravity given by the action
$S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(R+RF(\Box)R\Big)$ (28)
where $F(\Box)$ is an analytic function of the box operator, usually called the form factor BasiBeneito:2022wux , which we may expand as
$F(\Box)=\sum_{m=1}^{\infty}\frac{a_{m}}{M^{2m}}\Box^{m}=\sum_{m=1}^{\infty}f_{m}\Box^{m}\,.$ (29)
In the first representation $M^{2}$ carries a squared mass scale and consequently the coefficients $a_{m}$ are dimensionless constants. In the second form, the mass dependence is absorbed in $f_{m}$. In order to have more simplified expressions, we shall use the latter representation. Non-local theories of this kind, namely consisting of positive powers of the box operator are relevant in the UV regime Buoninfante:2022trw . In the metric formalism, it was shown in Biswas:2005qr that the class (48) gives rise to ghost and asymptotically free theories. Below we shall analyze the dynamics of these theories in the metric-affine approach.
Let us firstly derive the field equations associated with (48). Since this is the first time that non-local theories are studied in the Metric-Affine Gravity (MAG) framework, we will carefully provide some technical details that are needed for the derivation of the field equations. We will start with the metric variations which are the most elaborate. We start with the standard and trivial results
$\delta_{g}\sqrt{-g}=-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\delta g^{\mu\nu}$ (30)
$\delta_{g}R=(\delta g^{\mu\nu})R_{(\mu\nu)}$ (31)
Next, as a first step in dealing with the $\sqrt{-g}R(\delta\Box^{m})R$ term, that will appear after expanding $F$, we recall the identity
$\delta(\Box)\phi=\delta(\Box\phi)-\Box(\delta\phi)$ (32)
for any variation-$\delta$ and for an arbitrary scalar field $\phi$. Given that $\Box\phi=g^{\mu\nu}\widetilde{\nabla}_{\mu}\widetilde{\nabla}_{\nu}$, it immediately follows that
$\delta_{g}(\Box)\phi=(\delta g^{\mu\nu})\widetilde{\nabla}_{\mu}\widetilde{\nabla}_{\nu}\phi-g^{\mu\nu}(\delta_{g}\tilde{\Gamma}^{\lambda}{}_{\mu\nu})\partial_{\lambda}\phi$ (33)
Then, varying the Levi-Civita connection, we readily obtain
$\delta_{g}\tilde{\Gamma}^{\lambda}{}_{\mu\nu}=\frac{1}{2}g^{\alpha\lambda}\Big(\widetilde{\nabla}_{\mu}(\delta g_{\nu\alpha})+\widetilde{\nabla}_{\nu}(\delta g_{\mu\alpha})-\widetilde{\nabla}_{\alpha}(\delta g_{\mu\nu})\Big)$ (34)
Using the above two, after some partial integrations and using the identity $\delta g_{\alpha\beta}=-g_{\mu\alpha}g_{\nu\beta}\delta g^{\mu\nu}$, we find
$\sqrt{-g}f(\delta_{g}\Box)\phi=\sqrt{-g}(\delta g^{\mu\nu})\left(-(\widetilde{\nabla}_{(\mu}f)(\widetilde{\nabla}_{\nu)}\phi)+\frac{f}{2}(\Box\phi)g_{\mu\nu}+\frac{1}{2}(\widetilde{\nabla}_{\alpha}f)(\widetilde{\nabla}^{\alpha}\phi)g_{\mu\nu}\right)+t.d.$ (35)
up to total derivative terms $t.d.$ which we shall be dropping from now on. Note also the cancellation of the double derivative term $f\widetilde{\nabla}_{\mu}\widetilde{\nabla}_{\nu}\phi$ from the final expression. Now we have everything needed to tackle the variation of the $\sqrt{-g}RF(\Box)R$ term. Expressing the form factor in the power-series mentioned above, we have
$\delta_{g}(\sqrt{-g}RF(\Box)R)=\sum_{m=1}^{\infty}f_{m}\delta_{g}(\sqrt{-g}R\Box^{m}R)$ (36)
we isolate
$\delta_{g}(\sqrt{-g}R\Box^{m}R)=-\frac{1}{2}\sqrt{-g}R\Box^{m}Rg_{\mu\nu}(\delta g^{\mu\nu})+\sqrt{-g}(\delta_{g}R)\Box^{m}R+\sqrt{-g}R(\delta\Box^{m})R+\sqrt{-g}R\Box^{m}(\delta R)$ (37)
then, using the fact that up to total derivatives $\sqrt{-g}f\Box\phi=\sqrt{-g}(\Box f)\phi$ if we apply this m-times to the last term on the right-hand side of the above, we see that it is equal to the second. Collecting them and using $\delta_{g}R=(\delta g^{\mu\nu})R_{(\mu\nu)}$ it follows that
$\delta_{g}(\sqrt{-g}R\Box^{m}R)=-\frac{1}{2}\sqrt{-g}R\Box^{m}Rg_{\mu\nu}(\delta g^{\mu\nu})+\sqrt{-g}2R_{(\mu\nu)}(\Box^{m}R)(\delta g^{\mu\nu})+\sqrt{-g}R(\delta\Box^{m})R\,.$ (38)
Now only the last variation remains. In order to compute it, we use the identity
$\delta\Box^{m}=\sum_{l=0}^{m-1}\Box^{l}(\delta\Box)\Box^{m-l-1}$ (39)
along with (35), to which we set $f=\Box^{l}R=R^{[l]}$ and $\phi=\Box^{m-l-1}R=R^{[m-l-1]}$, to arrive at
$\sqrt{-g}R(\delta\Box^{m})R=\sum_{l=0}^{m-1}\sqrt{-g}R\Box^{l}(\delta_{g}\Box)\Box^{m-l-1}R=\sum_{l=0}^{m-1}\sqrt{-g}(\Box^{l}R)(\delta_{g}\Box)\Box^{m-l-1}R=$
$=\sqrt{-g}(\delta g^{\mu\nu})\sum_{l=0}^{m-1}\left(-(\widetilde{\nabla}_{(\mu}R^{[l]})(\widetilde{\nabla}_{\nu)}R^{[m-l-1]})+\frac{1}{2}R^{[l]}R^{[m-l]}g_{\mu\nu}+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R^{[l]})(\widetilde{\nabla}^{\alpha}R^{[m-l-1]})g_{\mu\nu}\right)$ (40)
Substituting the latter into (38), we obtain
$\delta_{g}(\sqrt{-g}R\Box^{m}R)=\sqrt{-g}(\delta g^{\mu\nu})\Big[-\frac{1}{2}R(\Box^{m}R)g_{\mu\nu}+2R_{(\mu\nu)}\Box^{m}R$
$+\sum_{l=0}^{m-1}\Big(-(\widetilde{\nabla}_{(\mu}R^{[l]})(\widetilde{\nabla}_{\nu)}R^{[m-l-1]})+\frac{1}{2}R^{[l]}R^{[m-l]}g_{\mu\nu}+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R^{[l]})(\widetilde{\nabla}^{\alpha}R^{[m-l-1]})g_{\mu\nu}\Big)\Big]$ (41)
Taking the $l=0$ value separately and consequently starting the sum from $l=1$ the first term in the parenthesis is canceled and we arrive at the final form
$\delta_{g}(\sqrt{-g}R\Box^{m}R)=\sqrt{-g}(\delta g^{\mu\nu})\Big[2R_{(\mu\nu)}\Box^{m}R-(\widetilde{\nabla}_{(\mu}R)(\widetilde{\nabla}_{\nu)}R^{[m-1]})+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R)(\widetilde{\nabla}^{\alpha}R^{[m-1]})g_{\mu\nu}$
$+\sum_{l=1}^{m-1}\Big(-(\widetilde{\nabla}_{(\mu}R^{[l]})(\widetilde{\nabla}_{\nu)}R^{[m-l-1]})+\frac{1}{2}R^{[l]}R^{[m-l]}g_{\mu\nu}+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R^{[l]})(\widetilde{\nabla}^{\alpha}R^{[m-l-1]})g_{\mu\nu}\Big)\Big]$ (42)
With this we are now in a position to derive the metric field equations corresponding to (11), which, after using all the above results, read
$R_{(\mu\nu)}-\frac{R}{2}g_{\mu\nu}+2R_{(\mu\nu)}F(\Box)R-(\widetilde{\nabla}_{(\mu}R)(\widetilde{\nabla}_{\nu)}F(\Box)\Box^{-1}R)+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R)(\widetilde{\nabla}^{\alpha}F(\Box)\Box^{-1}R)g_{\mu\nu}$
$+\sum_{m=2}^{\infty}\sum_{l=1}^{m-1}f_{m}\Big(-(\widetilde{\nabla}_{(\mu}R^{[l]})(\widetilde{\nabla}_{\nu)}R^{[m-l-1]})+\frac{1}{2}R^{[l]}R^{[m-l]}g_{\mu\nu}+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R^{[l]})(\widetilde{\nabla}^{\alpha}R^{[m-l-1]})g_{\mu\nu}\Big)\Big]=0$ (43)
where $\Box^{-1}$ is the inverse operator of $\Box$, i.e. $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$ and therefore $F(\Box)\Box^{-1}=\sum_{m=1}^{\infty}f_{m}\Box^{m-1}=f_{1}+f_{2}\Box+$... namely the sum start from a constant term. As a crosscheck, for $f_{1}=\gamma$, $f_{m}=0$, $m\geq 2$ it can be easily seen that the latter field equations are consistent with (13). Let us now proceed with the connection field equations. These are quite trivial since
$\delta_{\Gamma}(\sqrt{-g}R\Box^{m}R)=\sqrt{-g}(\delta_{\Gamma}R)\Box^{m}R+\sqrt{-g}R\Box^{m}(\delta_{\Gamma}R)=2\sqrt{-g}(\delta_{\Gamma}R)\Box^{m}R+t.d.$ (44)
where in the last equality we used the fact that $\sqrt{-g}f\Box^{m}\phi=\sqrt{-g}\phi\Box^{m}f+t.d.$, with $t.d.$ denoting total derivative terms. Then, the connection field equations are found to be
$\phi P_{\lambda}{}^{\mu\nu}+\delta_{\lambda}^{\nu}\partial^{\mu}\phi-g^{\mu\nu}\partial_{\lambda}\phi=0$ (45)
where $P_{\lambda}{}^{\mu\nu}$ is the Palatini tensor defined in (15) and
$\phi=1+2F(\Box)R$ (46)
Note that these are formally the same with (14), with the mere replacement $\gamma\Box R\rightarrow F(\Box)R$ and of course setting $\alpha=1$ and $\beta=0$. It is therefore again possible to solve for the connection. Indeed, despite the complexity of the metric field equations, the connection field equations obtained above, have a rather simple structure. As a matter of fact, since the form of (45) is identical to that of (14) with only the exact form of $\phi$ altered, the exact and unique solution to (45) is still
$N_{\alpha\mu\nu}=\frac{2}{(n-2)\phi}g_{\nu[\alpha}\partial_{\mu]}\phi+\frac{1}{2}g_{\alpha\mu}q_{\nu}$ (47)
with $\phi$ given by eq. (53). One can then proceed in a similar way to that of the previous section and bring the metric field equations in the form of Scalar-Tensor theories with higher order derivative terms for the scalar field.
IV.2 Non-local Infrared $\mathcal{L}=R+RF(\Box^{-1})R$ Theories
The previous results can be straightforwardly applied to non-local theories involving the inverse box operator Deser:2007jk $\Box^{-1}$ defined through $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$. As a prototype let us consider the class of non-local theories given by [^3] For the study of such theories in the metric formalism see e.g. Capozziello:2020xem ; Capozziello:2021bki ; Capozziello:2023mju .
$S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(R+RF(\Box^{-1})R\Big)$ (48)
where $F(\Box^{-1})$ is now an analytic function of the inverse box operator, which is expanded as
$F(\Box^{-1})=\sum_{m=1}^{\infty}c_{m}\Box^{-m}$ (49)
where $\Box^{-m}$ is an abbreviation for $(\Box^{-1})^{m}$. Note that this form of series for $F$, namely containing powers of the inverse box operator are relevant for the IR regime. In order to proceed with the field equations, we only need to vary the defining relation of $\Box^{-1}$ and subsequently multiply from the left by the same operator, to trivially get
$\delta\Box^{-1}=-\Box^{-1}(\delta\Box)\Box^{-1}$ (50)
With this result and using those of the previous subsection, varying with respect to the metric and the connection, we obtain the field equations
$R_{(\mu\nu)}-\frac{R}{2}g_{\mu\nu}+2R_{(\mu\nu)}F(\Box^{-1})R-\frac{g_{\mu\nu}}{2}RF(\Box^{-1})R$ (51)
$+\sum_{m=1}^{\infty}\sum_{l=0}^{m-1}c_{m}\Big(-(\widetilde{\nabla}_{(\mu}R^{-[l+1]})(\widetilde{\nabla}_{\nu)}R^{-[m-l]})+\frac{1}{2}R^{-[l+1]}R^{-[m-l-1]}g_{\mu\nu}+\frac{1}{2}(\widetilde{\nabla}_{\alpha}R^{-[l+1]})(\widetilde{\nabla}^{\alpha}R^{-[m-l]})g_{\mu\nu}\Big)\Big]=0$ (52)
and
$\Omega P_{\lambda}{}^{\mu\nu}+\delta_{\lambda}^{\nu}\partial^{\mu}\Omega-g^{\mu\nu}\partial_{\lambda}\Omega=0\;\;,\;\;\Omega=1+2F(\Box^{-1})R$ (53)
respectively. In the above we have used the notation $R^{-[m]}:=\Box^{-m}R=(\Box^{-1})^{m}R$. It is worth stressing again that the connection field equations are formally the same with (14), with $\phi$ replaced by $\Omega$. Finally, it goes without saying that one can combine the above results and derive the field equations for theories that contain both the box operator and its inverse at the same time. Let us now move on and develop some novel Non-local Metric-Affine Gravity theories which have no Riemannian analogue.
V Novel Theories with no Riemannian analogue
As we have already pointed out, the action (11) as well as its extension (48), have been well studied in the Riemannian case. In particular, in a Riemannian setting, (11) is the most general theory that is quadratic in scalar curvature. We note, however, that in MAG, squared torsion and non-metricity terms supplement the linear in curvature Einstein-Hilbert term. Therefore, in a MAG scheme before extending to (11) it would be quite natural to consider non-local actions consisting of scalars formed by sandwiching the box operator with torsion and non-metricity before moving to quadratic curvature. Given that there are 11 parity even quadratic invariants, the analogue of the theory (11) to MAG is
$S_{[}g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big[R+b_{1}S_{\alpha\mu\nu}S^{\alpha\mu\nu}+b_{2}S_{\alpha\mu\nu}S^{\mu\nu\alpha}+b_{3}S_{\mu}S^{\mu}+c_{1}Q_{\alpha\mu\nu}S^{\alpha\mu\nu}+c_{2}Q_{\mu}S^{\mu}+c_{3}q_{\mu}S^{\mu}$
$a_{1}Q_{\alpha\mu\nu}Q^{\alpha\mu\nu}+a_{2}Q_{\alpha\mu\nu}Q^{\mu\nu\alpha}+a_{3}Q_{\mu}Q^{\mu}+a_{4}q_{\mu}q^{\mu}+a_{5}Q_{\mu}q^{\mu}$
$b^{\prime}_{1}S_{\alpha\mu\nu}\Box S^{\alpha\mu\nu}+b^{\prime}_{2}S_{\alpha\mu\nu}\Box S^{\mu\nu\alpha}+b^{\prime}_{3}S_{\mu}\Box S^{\mu}+c^{\prime}_{1}Q_{\alpha\mu\nu}\Box S^{\alpha\mu\nu}+c^{\prime}_{2}Q_{\mu}\Box S^{\mu}+c^{\prime}_{3}q_{\mu}\Box S^{\mu}$
$a^{\prime}_{1}Q_{\alpha\mu\nu}\Box Q^{\alpha\mu\nu}+a^{\prime}_{2}Q_{\alpha\mu\nu}\Box Q^{\mu\nu\alpha}+a^{\prime}_{3}Q_{\mu}\Box Q^{\mu}+a^{\prime}_{4}q_{\mu}\Box q^{\mu}+a^{\prime}_{5}Q_{\mu}\Box q^{\mu}\Big]$ (54)
Some comments are in order. The first two lines consist of the usual Einstein-Hilbert term along with the quadratic invariants in torsion and non-metricity. These exhaust the number of parity even invariants up to dimension $L^{-2}]$. When only the first two lines of the above action are considered then the vacuum theory reduces to General Relativity Iosifidis:2021tvx . The third and fourth lines are the new non-local terms we consider. These have not been studied previously in the literature. Let qualitatively study their dynamics. The full field equations are quite lengthy and are given in the appendix. Suppressing indices, the connection field equations schematically look like
$\Box N\propto N$ (55)
where $N$ is the distortion. This shows that torsion and non-metricity propagate in this case and the theory has more degrees of freedom than General Relativity. It goes beyond the scope of the current study to investigate the healthy sub-sectors of this class of theories. However, it is worth mentioning that the non-local terms of the third and fourth lines, when added to the quadratic curvature MAG theories, may help to extend the healthy spectrum of the quadratic MAG. In order to get our point across, let us study a simple, but yet non-trivial, subsector of the above class.
V.1 An Example
Let us study the model given by
$S_{[}g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big[R+\alpha S_{\mu}S^{\mu}+\beta S_{\mu}\Box S^{\mu}\Big]$ (56)
which corresponds to setting all constants in (54) to zero apart from $b_{3}=\alpha$ and $b_{3}^{\prime}=\beta$. Varying with respect to the connection, we obtain
$P_{\lambda}{}^{\mu\nu}+2(\alpha S^{[\mu}+\beta\Box S^{[\mu})\delta^{\nu]}_{\lambda}=0$ (57)
Tracing in $\lambda=\mu$ and using the fact that $P_{\mu}{}^{\mu\nu}=0$, it follows that
$\alpha S_{\mu}+\beta\Box S_{\mu}=0$ (58)
which is a dynamical equation for the torsion vector. Therefore, in this case, torsion propagates in vacuum due to the inclusion of the non-local term. Substituting this back to (57) gives $P_{\lambda}{}^{\mu\nu}=0$ and consequently the distortion is readily computed to be
$N_{\alpha\mu\nu}=\frac{1}{2}g_{\alpha\mu}q_{\nu}\,.$ (59)
With this result, we then find the full forms of torsion and non-metricity
$S_{\mu\nu\alpha}=\frac{2}{(n-1)}S_{[\mu}g_{\nu]\alpha}\;\;,\;\;Q_{\alpha\mu\nu}=\frac{Q_{\alpha}}{n}g_{\mu\nu}$ (60)
where the vectors are related through
$Q_{\mu}=nq_{\mu}=-\frac{4n}{(n-1)}S_{\mu}$ (61)
and the whole dynamics is contained in (58). It has to be noted that the distortion (59) is of a very peculiar type (purely projective mode). It has the remarkable property to give no post-Riemannian contributions to the symmetric part of the Ricci tensor and, consequently, also the Ricci scalar, is
$R_{(\mu\nu)}=\tilde{R}_{\mu\nu}\;\;,\;\;R=\tilde{R}\,$. (62)
This is of course tightly related with the fact that two connections are related through a projective transformation and have their symmetric Ricci tensor equal. Furthermore the connections, in this equivalence class, share also the same autoparallel curves up to a reparameterization of the affine parameter Iosifidis:2018zjj .
Finally, varying with respect to the connection and using the form of (59) we find the metric field equations
$\tilde{R}_{\mu\nu}-\frac{1}{2}\tilde{R}g_{\mu\nu}=\frac{\alpha}{2}S_{\alpha}S^{\alpha}g_{\mu\nu}-\alpha S_{\mu}S_{\nu}-\frac{\beta}{2}(\widetilde{\nabla}_{\alpha}S^{\alpha})^{2}g_{\mu\nu}$
$+\beta\Big(\widetilde{\nabla}_{\mu}S_{\alpha}\widetilde{\nabla}_{\nu}S^{\alpha}+\widetilde{\nabla}_{\alpha}S_{\mu}\widetilde{\nabla}^{\alpha}S_{\nu}\Big)-\frac{\beta}{2}\widetilde{\nabla}^{\alpha}\Big(S_{(\mu}\widetilde{\nabla}_{\nu)}S_{\alpha}+(\widetilde{\nabla}_{\alpha}S_{(\mu})S_{\nu)}-S_{\alpha}\widetilde{\nabla}_{(\mu}S_{\nu)}\Big)$ (63)
from which it is clear that the resulting theory is a Vector-Tensor one, where the vector is of geometric origin. Of course, it goes without saying that in general, these excitations extra degrees of freedom will lead to pathologies. However, it should be noted that the additional terms we consider here when added to the quadratic MAG theory could extend the spectrum of healthy theories just like the inclusions of cubic terms in the action, see Bahamonde:2024efl .
VI Applications to Cosmology
Let us now discuss some applications for our development. Consider a flat Fiedmann-Lemaìtre-Roberston-Walker (FLRW) Universe with the usual Robertson-Walker line element in Cartesian coordinates $ds^{2}=-dt^{2}+a^{2}(t)(dx^{2}+dy^{2}+dz^{2})$, where $a(t)$ is the scale factor. We shall first derive the modified Friedmann equations for the model (11). Taking the $00$ and $ij$ components of the resulting field equations (27), we find
$3H^{2}=-\frac{1}{4}\left(\frac{\dot{\phi}}{\phi}\right)^{2}-3H\left(\frac{\dot{\phi}}{\phi}\right)+\frac{1}{\phi}\left(\gamma\dot{\psi}^{2}+\frac{\psi}{2}\right)\,$, (64)
$\frac{\ddot{a}}{a}=\frac{5}{12}\left(\frac{\dot{\phi}}{\phi}\right)^{2}-\frac{1}{3}H\left(\frac{\dot{\phi}}{\phi}\right)-\frac{\ddot{\phi}}{2\phi}+\frac{1}{6\phi}(\psi-\gamma\dot{\psi}^{2})\,$, (65)
where $\phi=1-6\gamma(\ddot{\psi}+3H\dot{\psi})$. In addition, the evolution eq. (25) for the scalar field takes the form
$2\gamma\psi(\ddot{\psi}+3H\dot{\psi})+\psi+\gamma\dot{\psi}^{2}=0$ (66)
The latter together with the first Friedmann equation above, fully determine the dynamics of the system. Note that for $\gamma=0$, it follows that $\psi=0$ and $\phi=1$ and the above reduce to the usual Friedmann equations as expected. Let us find solutions to the above system. With the premise that non-local effects (attributed to the scalar $\phi$) manifest themselves at the early stages of the Universe history, it is meaningful to consider the ansatz
$\psi=\psi_{0}e^{-bt}$ (67)
with $b>0$. The value of this parameter is obtained by imposing consistency on the system of equations. The above ansatz ensures that non-local quantum-mechanical effects are important at very early times and quickly diminish as we enter into the classical regime. Substituting (67) into (66), it then follows that
$H=\lambda be^{bt}+\frac{b}{2}$ (68)
where $\lambda=\frac{1}{6\gamma b^{2}\psi_{0}}$. The latter immediately integrates to
$a(t)=a_{0}e^{{\lambda}e^{bt}+\frac{b}{2}t}$ (69)
namely, non-locality produces a superexponential expansion. Further substitution of (67) and (68) fixes the value of b. This simple application demonstrates the importance of early non-local effects in non-Riemannian Universes.
For completeness, we will also report the corresponding cosmological equations for the novel theory (56). Let us recall that, in this case, since only the vector part of torsion is excited, in the FLRW spacetime the full torsion is fully described by a single scalar field and the torsion vector takes the form tsamparlis1979cosmological
$S_{\mu}=3\Phi u_{\mu}$ (70)
Now, contracting (58) with $S^{\mu}$ and using the identity $S^{\mu}\Box S_{\mu}=\frac{1}{2}\Box(S_{\mu}S^{\mu})-\tilde{\nabla}^{\alpha}S^{\beta}\tilde{\nabla}_{\alpha}S_{\beta}$ along with the above form of torsion, we readily obtain the evolution equation
$\alpha\Phi^{2}+\frac{\beta}{6}\Big(\ddot{(\Phi^{2})}+3H\dot{(\Phi)^{2}}\Big)+\beta(\dot{\Phi})^{2}+3\beta H^{2}\Phi^{2}=0$ (71)
for the torsion field. Finally, taking the trace of (63), we find a variant of the Friedmann equation
$2\left(H^{2}+\frac{\ddot{a}}{a}\right)=-3\alpha\Phi^{2}+\frac{9\beta}{2}(\dot{\Phi}+3H\Phi)^{2}-6\beta(\dot{\Phi}+3H^{2}\Phi^{2})+3\beta(\ddot{\Phi}\Phi+\dot{\Phi}^{2}+3H\Phi\dot{\Phi})-\frac{\beta}{2}\Phi(\ddot{\Phi}+3\dot{H}\Phi+3H\dot{\Phi})$ (72)
The latter two equations fully describe the homogeneous cosmological dynamics of the model (56). It goes beyond the purpose of the current work to further analyze such cosmological models, however let us note, as seen from these examples, that non-local MAG offers a rich cosmological dynamics.
VII Discussion and Conclusions
We have constructed the metric-affine version of Non-local Gravity. Starting with a simple non-local metric-affine model (see eq. (11)), after solving the associated connection field equations, we showed that the theory is on-shell equivalent to a metric Scalar-Tensor Gravity with higher order derivative terms for the propagating scalar field. We then generalized the results to generic Infinite Derivative Gravity theories of the type $\mathcal{L}=RF(\Box)R$, where $R$ is the general scalar curvature and $F(\Box)$ is an analytic function of the box operator. In particular, after devising the general method to obtain the field equations, we again successfully solved the connection field equations and obtained an on-shell equivalence to some higher derivative Scalar-Tensor Gravity. The same analysis was also performed for non-local theories that are relevant in the IR regime, namely of the form $\mathcal{L}=RF(\Box)R$ where $F(\Box^{-1})$ is now an analytic function of the inverse box operator $\Box^{-1}$, defined through $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$. It was shown that also in this case the connection field equations can be solved analytically and an equivalence to higher derivative Scalar-Tensor Theories was established in this case as well.
Furthermore, we constructed a novel non-local theory, as given by the action (54), which has no Riemannian analogue. In particular, our proposed theory consists of non-local quantum modifications to torsion and non-metricity that supplement the usual quadratic (in torsion and non-metricity) MAG. For this novel construction, we discussed how both torsion and non-metricity become dynamical even in vacuum. In general, when all modes of torsion and non-metricity are excited, the theory is equivalent to some generic Scalar-Vector-Tensor theory. Consequently, in this setting, such generalizations, already existing in the literature and including additional scalar and vector fields, here arise naturally from the enrichment of the underlying geometry. In this sense, the extra fields have a geometric origin, coming from torsion and non-metricity. Finally, we briefly discussed possible cosmological applications of non-local MAG. The full cosmological analysis of such theories will be studied elsewhere.
Acknowledgements
The authors acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), Sezioni di Napoli e di Torino, Iniziative Specifiche QGSKY and MOONLIGHT2. This paper is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).
VIII Appendix
VIII.1 Non-Riemannian expansions
Let us gather here some useful formulae that were used in the various derivations. Firstly, in the vanishing non-metricity gauge, namely for the distortion
$N_{\alpha\mu\nu}=\frac{2}{(n-2)}g_{\nu[\alpha}u_{\mu]}$ (73)
where we have set
$u_{\alpha}:=\frac{\partial_{\alpha}\phi}{\phi}=\partial_{\alpha}\ln{|\phi|}\;,$ (74)
we compute the non-Riemannian expansions
$R_{(\mu\nu)}=\tilde{R}_{\mu\nu}-\tilde{\nabla}_{(\mu}u_{\nu)}+\frac{1}{(n-2)}u_{\mu}u_{\nu}-\frac{1}{(n-2)}\Big(\tilde{\nabla}_{\alpha}u^{\alpha}+u_{\alpha}u^{\alpha}\Big)g_{\mu\nu}$ (75)
and
$R=\tilde{R}-\left(\frac{n-1}{n-2}\right)\Big(2\tilde{\nabla}_{\alpha}u^{\alpha}+u_{\alpha}u^{\alpha}\Big)$ (76)
From which it follows that
$R_{(\mu\nu)}-\frac{R}{2}g_{\mu\nu}=\tilde{R}_{\mu\nu}-\frac{\tilde{R}}{2}g_{\mu\nu}-\tilde{\nabla}_{(\mu}u_{\nu)}+\left[\tilde{\nabla}_{\alpha}u^{\alpha}+\frac{(n-3)}{2(n-2)}u_{\alpha}u^{\alpha}\right]g_{\mu\nu}$ (77)
VIII.2 Variations
Here we provide some details related to the variations, since this is the first time this framework is considered. Let us denote with $\delta_{\chi}$ the variation with respect to the $\chi$ which in our case would be either the metric $g$ or the connection $\Gamma$. The first thing to note is that since our box operator contains only the metric, its $\Gamma$-variation immediately vanishes, viz.
$\delta_{\Gamma}\Box=0$ (78)
With this fact let us proceed to derive the connection field equations for the Theory (54). The variations for the first two lines in (54) are well known, see e.g. Iosifidis:2021tvx , so we will concentrate on the derivation of the new part. Set
$\mathcal{L}_{\Box}^{(2)}:=b^{\prime}_{1}S_{\alpha\mu\nu}\Box S^{\alpha\mu\nu}+b^{\prime}_{2}S_{\alpha\mu\nu}\Box S^{\mu\nu\alpha}+b^{\prime}_{3}S_{\mu}\Box S^{\mu}+c^{\prime}_{1}Q_{\alpha\mu\nu}\Box S^{\alpha\mu\nu}+c^{\prime}_{2}Q_{\mu}\Box S^{\mu}+c^{\prime}_{3}q_{\mu}\Box S^{\mu}$
$a^{\prime}_{1}Q_{\alpha\mu\nu}\Box Q^{\alpha\mu\nu}+a^{\prime}_{2}Q_{\alpha\mu\nu}\Box Q^{\mu\nu\alpha}+a^{\prime}_{3}Q_{\mu}\Box Q^{\mu}+a^{\prime}_{4}q_{\mu}\Box q^{\mu}+a^{\prime}_{5}Q_{\mu}\Box q^{\mu}$ (79)
Then after some partial integrations and relabeling of the indices we can gather
$\delta_{\Gamma}\Big(\sqrt{-g}\mathcal{L}_{\Box}^{(2)}\Big)=(\Box\Sigma^{\alpha\mu\nu})\delta_{\Gamma}S_{\alpha\mu\nu}+(\Box\Omega^{\alpha\mu\nu})\delta_{\Gamma}Q_{\alpha\mu\nu}+t.d.$ (80)
where
$\Sigma^{\alpha\mu\nu}=2b_{1}^{\prime}S^{\alpha\mu\nu}+b_{2}^{\prime}S^{\nu\alpha\mu}+b_{2}^{\prime}S^{\mu\nu\alpha}+2b_{3}^{\prime}S^{\alpha}g^{\mu\nu}+c_{1}^{\prime}Q^{\alpha\mu\nu}+c_{2}^{\prime}Q^{\alpha}g^{\mu\nu}+c_{3}^{\prime}q^{\alpha}g^{\mu\nu}$ (81)
$\Omega^{\alpha\mu\nu}=2a_{1}^{\prime}Q^{\alpha\mu\nu}+a_{2}^{\prime}Q^{\nu\alpha\mu}+a_{2}^{\prime}Q^{\mu\nu\alpha}+c_{1}^{\prime}S^{\alpha\mu\nu}+2a_{3}^{\prime}Q^{\alpha}g^{\mu\nu}+2a_{4}^{\prime}q^{\nu}g^{\alpha\mu}+a_{5}^{\prime}Q^{\nu}g^{\alpha\mu}+a_{5}^{\prime}q^{\alpha}g^{\mu\nu}+c_{2}^{\prime}Q^{\alpha}g^{\mu\nu}+c_{3}^{\prime}S^{\nu}g^{\alpha\mu}$ (82)
Then using the variations $\delta_{\Gamma}S_{\alpha\beta}{}{}^{\lambda}=\delta^{[\mu}_{\alpha}\delta^{\nu]}_{\beta}\delta\Gamma^{\lambda}{}_{\mu\nu}$ and $\delta_{\Gamma}Q_{\rho\alpha\beta}=2\delta^{\nu}_{\rho}\delta^{\mu}_{(\alpha}g_{\beta)\lambda}\delta\Gamma^{\lambda}{}_{\mu\nu}$ we finally arrive at
$\delta_{\Gamma}\Big(\sqrt{-g}\mathcal{L}_{\Box}^{(2)}\Big)=(\delta\Gamma^{\lambda}{}_{\mu\nu})\Big(\Box(\Sigma^{[\mu\nu]}{}{}_{\lambda}+2\Omega^{\nu(\mu\alpha)}g_{\alpha\lambda})\Big)$ (83)
Consequently, using also known results for the $\Gamma$-variation of the rest of the terms in (54), we find the connection field equations
$\left(\frac{Q_{\lambda}}{2}+2S_{\lambda}\right)g^{\mu\nu}-Q_{\lambda}^{\;\;\mu\nu}-2S_{\lambda}^{\;\;\mu\nu}+\left(q^{\mu}-\frac{Q^{\mu}}{2}-2S^{\mu}\right)\delta_{\lambda}^{\nu}+4a_{1}Q^{\nu\mu}_{\;\;\;\;\lambda}+2a_{2}(Q^{\mu\nu}_{\;\;\;\;\lambda}+Q_{\lambda}^{\;\;\;\mu\nu})+2b_{1}S^{\mu\nu}_{\;\;\;\;\lambda}$
$+2b_{2}S_{\lambda}^{\;\;\;[\mu\nu]}+c_{1}\Big(S^{\nu\mu}_{\;\;\;\;\lambda}-S_{\lambda}^{\;\;\;\nu\mu}+Q^{[\mu\nu]}_{\;\;\;\;\;\lambda}\Big)+\delta_{\lambda}^{\mu}\Big(4a_{3}Q^{\nu}+2a_{5}q^{\nu}+2c_{2}S^{\nu}\Big)+\delta_{\lambda}^{\nu}\Big(a_{5}Q^{\mu}+2a_{4}q^{\mu}+c_{3}S^{\mu}\Big)$
$+g^{\mu\nu}\Big(a_{5}Q_{\lambda}+2a_{4}q_{\lambda}+c_{3}S_{\lambda}\Big)+\Big(c_{2}Q^{[\mu}+c_{3}q^{[\mu}+2b_{3}S^{[\mu}\Big)\delta^{\nu]}_{\lambda}+\Big(\Box(\Sigma^{[\mu\nu]}{}{}_{\lambda}+2\Omega^{\nu(\mu\alpha)}g_{\alpha\lambda})\Big)=0$ (84)
From this and using the relations $S_{\mu\nu\alpha}=N_{\alpha[\mu\nu]}$ and $Q_{\alpha\mu\nu}=2N_{(\mu\nu)\alpha}$ it is then clear that we have a dynamical equation for the distortion which we schematically express as $\Box N=N$ as stated.Let us now illustrate how one can derive the metric field equations for the action (11). The $g$-variation first two lines that contain no box operator is well known, see e.g. Iosifidis:2021bad . The lines containing the box operator are trickier so let us elaborate on them in detail. We shall perform the variation on the first term $\sqrt{-g}S_{\mu\nu\alpha}\Box S^{\mu\nu\alpha}$ and of course the same procedure must then be followed for the rest of the terms. The first thing to do is to perform a partial integration
$\sqrt{-g}S_{\alpha\mu\nu}\Box S^{\alpha\mu\nu}=\sqrt{-g}g^{\kappa\lambda}\tilde{\nabla}_{\kappa}S_{\alpha\mu\nu}\tilde{\nabla}_{\lambda}S^{\alpha\mu\nu}+t.d.$ (85)
then extract all g-dependence from the mixed torsion
$\sqrt{-g}S_{\alpha\mu\nu}\Box S^{\alpha\mu\nu}=\sqrt{-g}g^{\kappa\lambda}g_{\gamma\rho}g^{\alpha\delta}g^{\beta\epsilon}\tilde{\nabla}_{\kappa}S_{\alpha\beta}{}{}^{\rho}\tilde{\nabla}_{\lambda}S_{\delta\epsilon}{}{}^{\gamma}+t.d.$ (86)
in order to use the fact that the prototype torsion $S_{\mu\nu}{}{}^{\rho}$ is independent from the metric. Employing the Leibniz rule, we obtain
$\delta_{g}(\sqrt{-g}S_{\alpha\mu\nu}\Box S^{\alpha\mu\nu})=\sqrt{-g}(\delta g^{\mu\nu})\Big[-\frac{1}{2}g_{\mu\nu}(\tilde{\nabla}_{\lambda}S_{\alpha\beta\gamma})(\tilde{\nabla}^{\lambda}S^{\alpha\beta\gamma})+(\tilde{\nabla}_{\mu}S_{\alpha\beta\gamma})(\tilde{\nabla}_{\nu}S^{\alpha\beta\gamma})$
[…]
We now isolate the last term. Since $\delta_{g}S_{\alpha\beta}{}{}^{\gamma}=0$ we have that
$\delta_{g}(\tilde{\nabla}_{\lambda}S_{\alpha\beta}{}{}^{\gamma})=-(\delta_{g}\tilde{\Gamma}^{\kappa}{}_{\alpha\lambda})S_{\kappa\beta}{}{}^{\gamma}-(\delta_{g}\tilde{\Gamma}^{\kappa}{}_{\beta\lambda})S_{\alpha\kappa}{}{}^{\gamma}+(\delta_{g}\tilde{\Gamma}^{\gamma}{}_{\kappa\lambda})S_{\alpha\beta}{}{}^{\kappa}$ (88)
Using this, after relabeling indices and gathering terms it follows that
[…]
and upon using (34) we finally find
$2\sqrt{-g}(\tilde{\nabla}^{\lambda}S^{\alpha\beta}{}{}_{\gamma})\delta_{g}(\tilde{\nabla}_{\lambda}S_{\alpha\beta}{}{}^{\gamma})=(\delta g^{\mu\nu})\sqrt{-g}\tilde{\nabla}^{\lambda}\Big(S_{\alpha\beta(\mu}\tilde{\nabla}_{\nu)}S^{\alpha\beta}{}{}_{\lambda}-2S_{(\mu}{}{}^{\alpha\beta}\tilde{\nabla}_{\nu)}S_{\lambda\alpha\beta}+S_{\alpha\beta(\mu}\tilde{\nabla}_{|\lambda|}S^{\alpha\beta}{}{}_{\nu)}$
[…]
Consequently, for the full variation we have
$\delta_{g}(\sqrt{-g}S_{\alpha\mu\nu}\Box S^{\alpha\mu\nu})=\sqrt{-g}(\delta g^{\mu\nu})\Big[-\frac{1}{2}g_{\mu\nu}(\tilde{\nabla}_{\lambda}S_{\alpha\beta\gamma})(\tilde{\nabla}^{\lambda}S^{\alpha\beta\gamma})+(\tilde{\nabla}_{\mu}S_{\alpha\beta\gamma})(\tilde{\nabla}_{\nu}S^{\alpha\beta\gamma})$
$-(\tilde{\nabla}^{\lambda}S_{\alpha\beta\nu})(\tilde{\nabla}^{\lambda}S^{\alpha\beta}{}{}_{\mu})+2(\tilde{\nabla}_{\lambda}S_{\mu\alpha\beta})(\tilde{\nabla}^{\lambda}S_{\nu}{}^{\alpha\beta})\tilde{\nabla}^{\lambda}\Big(S_{\alpha\beta(\mu}\tilde{\nabla}_{\nu)}S^{\alpha\beta}{}{}_{\lambda}-2S_{(\mu}{}{}^{\alpha\beta}\tilde{\nabla}_{\nu)}S_{\lambda\alpha\beta}+S_{\alpha\beta(\mu}\tilde{\nabla}_{|\lambda|}S^{\alpha\beta}{}{}_{\nu)}$
[…]
The same procedure is then followed for the rest of the terms of this kind.

### v2

@@
 These quantum corrections, that alter the classical gravitational action, encode non-locality and dramatically change the associated dynamics producing new degrees of freedom.
-Non-locality is typically expressed through the inclusion of inverse powers of the box operator $\Box^{-1}$ sandwiched between geometric quantities (like the Ricci scalar) that supplement the gravitational action and consequently the associated field equations Mashhoon:2017qyw ; Deser:2007jk ; Capozziello:2008gu ; Nojiri:2010pw .
+Non-locality is typically expressed through the inclusion of inverse powers of the box operator $\Box^{-1}$ sandwiched between geometric quantities (like the Ricci scalar) that supplement the gravitational action and consequently the associated field equations Mashhoon (2017); Deser and Woodard (2007); Capozziello et al.
+(2009); Nojiri et al.
+(2011).
 Another kind of non-locality is realized by adding an infinite series of the box operator in the form of entire functions like e.g.
 $RF(\Box)R$ where $R$ is the Ricci scalar and $F(\Box)$ a holomorphic function of the box operator.
-This is known as Infinite Derivative Gravity (IDG) efimov1967non .
-Besides fundamental physics issues, the key idea to adopt non-local gravity models is to obtain characteristic lengths and masses capable of addressing Ultra-Violet and Infra-Red issues in astrophysics and cosmology Blome:2010xn ; Roshan:2022zov .
+This is known as Infinite Derivative Gravity (IDG) Efimov (1967).
+Besides fundamental physics issues, the key idea to adopt non-local gravity models is to obtain characteristic lengths and masses capable of addressing Ultra-Violet and Infra-Red issues in astrophysics and cosmology Blome et al.
+(2010); Roshan and Mashhoon (2022); Boos (2020); Boos et al.
+(2018).
 Cosmological applications of non-local gravity have been studied in Refs.
-Deser:2007jk ; Koivisto:2008xfa ; Calcagni:2018pro ; Capozziello:2024qol ; Capozziello:2021krv ; Nojiri:2019dio ; Biswas:2010zk ; Kolar:2021qox ; Koshelev:2016xqb ; Calcagni:2013vra ; Dirian:2014bma and some problems like the Weinberg no-go theorem for cosmological constant can be addressed in this framework Capozziello:2025bsm .
-The teleparallel formulation of non-local gravity along with its cosmological applications, has been considered in Bahamonde:2017bps .
-Also black hole solutions can be improved with the presence of non-local terms Bambi:2016uda ; Capozziello:2025iwn as well as gravitational waves where new polarizations emerge Capozziello:2024bjb ; Foffa:2013sma .
+Deser and Woodard (2007); Koivisto (2008); Calcagni et al.
+(2018); Capozziello et al.
+(2024); Capozziello and Bajardi (2022); Nojiri et al.
+(2020); Biswas et al.
+(2010); Kolář et al.
+(2022); Koshelev et al.
+(2016); Calcagni et al.
+(2014); Dirian et al.
+(2015) and some problems like the Weinberg no-go theorem for cosmological constant can be addressed in this framework Capozziello et al.
+(2025c).
+Non-local hybrid metric-Palatini cosmology was studied in Bombacigno et al.
+(2025).
+The teleparallel formulation of non-local gravity along with its cosmological applications, has been considered in Bahamonde et al.
+(2017).
+Also black hole solutions can be improved with the presence of non-local terms Bambi et al.
+(2016); Capozziello et al.
+(2025d) as well as gravitational waves where new polarizations emerge Capozziello et al.
+(2025a); Foffa et al.
+(2014).
 Here we are going to formulate non-local gravity in the metric-affine geometry setup in order to consider the most general framework where these theories can be formulated.
-This approach proves to be useful also in view of extensions and modifications of General Relativity where the geometric invariants can be other than the standard curvature Bahamonde:2017bps ; BeltranJimenez:2019esp ; Capozziello:2022zzh ; Capozziello:2025hyw ; Golovnev:2023yla .
+This approach proves to be useful also in view of extensions and modifications of General Relativity where the geometric invariants can be other than the standard curvature Bahamonde et al.
+(2017); Beltrán Jiménez et al.
+(2019); Capozziello et al.
+(2022, 2025b); Golovnev (2023).
 The paper is organized as follows.
@@
 $III$, we start with the metric-affine analogue of the theory introduced in Ref.
-Gottlober:1989ww , which was studied in the Riemannian setting.
+Gottlober et al.
+(1990), which was studied in the Riemannian setting.
 Immediately the differences with the metric formalism emerge.
@@
 The difference between the general connection $\Gamma^{\lambda}{}_{\mu\nu}$ and the Levi-Civita one, $\tilde{\Gamma}^{\lambda}{}_{\mu\nu}$, defines a tensor that quantifies how much the geometry deviates from being Riemannian.
-This is the so-called distortion tensor schouten1954ricci ; hehl1995metric
+This is the so-called distortion tensor Schouten (1954); Hehl et al.
+(1995a)
 $N^{\lambda}{}_{\mu\nu}:=\Gamma^{\lambda}{}_{\mu\nu}-\tilde{\Gamma}^{\lambda}{}_{\mu\nu}$ (5)
@@
 Torsion and non-metricity can then be recovered from the distortion from the relations (see e.g.
-Hehl:1994ue ),
+Hehl et al.
+(1995b)),
 $Q_{\mu\alpha\beta}=2N_{(\alpha\beta)\mu}\;\;,\;\;S_{\mu\nu\alpha}=N_{\alpha[\mu\nu]}.$ (7)
@@
 In this form, it can further act upon tensor fields of arbitrary rank, $\Box T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}=\widetilde{\nabla}^{\mu}\widetilde{\nabla}_{\mu}T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}$.
-Note that in Briscese:2015urz the definition
+Note that in Briscese and Pucheu (2016) the definition
 $\Box^{(\Gamma)}=g^{\mu\nu}\nabla_{\mu}\nabla_{\nu},$ (10)
@@
 Indeed, operators like $\nabla_{\mu}(g^{\mu\nu}\nabla_{\nu})$, $\frac{1}{\sqrt{-g}}\nabla_{\mu}(\sqrt{-g}g^{\mu\nu}\nabla_{\nu})$ or $\nabla_{\mu}(\nabla_{\nu}g^{\mu\nu})$ also reduce to $\partial^{\mu}\partial_{\nu}$ in the flat space limit.
-One would then need to take all possible linear combinations of such operators (with the coefficients added up to one) in order not to discriminate among them, see for instance Shimada:2018lnm .
-However, this approach, not only does it complicate the analysis but also introduces new arbitrary parameters.
+One would then need to take all possible linear combinations of such operators (with the coefficients added up to one) in order not to discriminate among them, see for instance Shimada et al.
+(2019).
+However, this approach, not only does it complicate the analysis but also introduces new arbitrary parameters [^2] As an example, let us consider the linear combination $\Box=ag^{\mu\nu}\nabla_{\mu}\nabla_{\nu}+b\nabla_{\mu}g^{\mu\nu}\nabla_{\nu}$.
+Consistency in the flat space limit then demands that $a+b=1$.
+Furthermore using the connection decomposition (5) it follows that $\Box f=\Box_{g}f+\Big(\frac{Q^{\mu}}{2}+2S^{\mu}+(b-1)q^{\mu}\Big)\partial_{\mu}f$.
+That is, it can always be split into the Riemannian box plus derivative couplings of the function and the non-Riemannian vectors..
 Thirdly, for operators like (10), terms like $R_{\mu\nu}\Box^{(\Gamma)}R^{\mu\nu}$ and $R^{\mu\nu}\Box^{(\Gamma)}R_{\mu\nu}$ do not coincide since the metric cannot freely move through the full covariant derivative $\nabla_{\mu}$.
@@
 It is therefore only natural to stick to (8) for the definition of the box operator in our current context.
+[^3] The same definition of the box operator was also used in Bombacigno et al.
+(2025).
 III A simple Warm-up Theory
-Let us start with the model in Gottlober:1989ww , but now in the metric-affine formalism, where the connection and the metric are independent fields. We have
+Let us start with the model in Gottlober et al. (1990), but now in the metric-affine formalism, where the connection and the metric are independent fields. We have
 $S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(\alpha R+\beta R^{2}+\gamma R\Box R\Big)$ (11)
@@
 Let us also note that, in this action, the scalar curvature $R$ is now the generic one that has also torsion and non-metricity.
-The action (11) is, up to surface terms, equivalent to
+The action (11) is, up to surface terms, equivalent to [^5] Of course when acting all scalars, the covariant derivative, of any connection, always reduces to the partial one Schouten (1954).
+Therefore, for the scalar curvature we have $\nabla_{\mu}R=\widetilde{\nabla}_{\mu}R=\partial_{\mu}R$.
 $S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(\alpha R+\beta R^{2}-\gamma g^{\mu\nu}\partial_{\mu}R\partial_{\nu}R\Big)\,.$ (12)
@@
 $(\partial R)^{2}=g^{\mu\nu}\partial_{\mu}R\partial_{\nu}R\;\;,\;\;\phi=\alpha+2\beta R+2\gamma\Box R$ (16)
-Using the method developed in Iosifidis:2018jwu , from the above connection field equations, one finds the distortion
+Using the method developed in Iosifidis (2019), from the above connection field equations, one finds the distortion
 $N_{\alpha\mu\nu}=\frac{2}{(n-2)\phi}g_{\nu[\alpha}\partial_{\mu]}\phi+\frac{1}{2}g_{\alpha\mu}q_{\nu}$ (17)
-Now notice that our action (11) is invariant under projective transformations of the connection
+and consequently, torsion and non-metricity read
+$S_{\mu\nu}{}{}^{\lambda}=\delta^{\lambda}_{[\mu}\left(\frac{1}{2}q_{\nu]}-\frac{1}{(n-2)}\frac{\partial_{\nu]}\phi}{\phi}\right)$ (18)
+$Q_{\nu\alpha\mu}=\frac{1}{2}g_{\alpha\mu}q_{\nu}$ (19)
+Let us observe that both are of purely vectorial type. Now notice that our action (11) is invariant under projective transformations of the connection
 $\Gamma^{\lambda}{}_{\mu\nu}\mapsto\Gamma^{\lambda}{}_{\mu\nu}+\delta^{\lambda}_{\mu}\xi_{\nu}$ (20)
 since it is built solely from $R$ which is itself projectively invariant.
-This means that there is an undetermined vectorial degree of freedom which can be fixed by a gauge choice iosifidis2019exactly .
+This means that there is an undetermined vectorial degree of freedom which can be fixed by a gauge choice Iosifidis (2019) [^6] It should be noted that this invariance has implications when matter is added.
+In particular, it forces the dilation part of the hypermomentum to be vanishing (see for instance Hehl et al.
+(1981)).
+Of course this constraint is avoided if the matter does not couple to the connection but only to the metric..
 Two special gauge choices are of interest here.
@@
 This is the vanishing torsion gauge.
-We see therefore a torsion/non-metricity duality similar to the one that appears in $f(R)$ theories Iosifidis:2018zjj .
+We see therefore a torsion/non-metricity duality similar to the one that appears in $f(R)$ theories Iosifidis et al.
+(2019).
 Let us now investigate the metric field equations (13).
@@
 We see therefore that the initial theory is on-shell equivalent to a Scalar-Tensor theory with higher order derivatives for the scalar field.
-It is important to stress that, in constrast to the Riemannian case where the resulting metric field equations contain sixth order derivatives of the scalar field, in the metric-affine approach developed here, we have up to fourth-order derivatives of the scalar field.
-In order words, with respect to the metric formalism developed in Gottlober:1989ww , the metric-affine approach reduces the dynamics.
-See also Amendola:1993bg for a cosmological application.
+It is important to stress that, in contrast to the Riemannian case where the resulting metric field equations contain sixth order derivatives of the scalar field, in the metric-affine approach developed here, we have up to fourth-order derivatives of the scalar field.
+In order words, with respect to the metric formalism developed in Gottlober et al.
+(1990) [^7] See also Amendola et al.
+(1993) for a cosmological application., the metric-affine approach reduces the order of the derivatives describing the dynamics of the extra scalar field.
 IV Non-local Metric-Affine Theories
@@
 $S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(R+RF(\Box)R\Big)$ (30)
-where $F(\Box)$ is an analytic function of the box operator, usually called the form factor BasiBeneito:2022wux , which we may expand as
+where $F(\Box)$ is an analytic function of the box operator, usually called the form factor Bas i Beneito et al. (2024), which we may expand as
 $F(\Box)=\sum_{m=1}^{\infty}\frac{a_{m}}{M^{2m}}\Box^{m}=\sum_{m=1}^{\infty}f_{m}\Box^{m}\,.$ (31)
@@
 In order to have more simplified expressions, we shall use the latter representation.
-Non-local theories of this kind, namely consisting of positive powers of the box operator are relevant in the UV regime Buoninfante:2022trw .
-In the metric formalism, it was shown in Biswas:2005qr that the class (48) gives rise to ghost and asymptotically free theories.
+Non-local theories of this kind, namely consisting of positive powers of the box operator are relevant in the UV regime Buoninfante et al.
+(2022).
+In the metric formalism, it was shown in Biswas et al.
+(2006) that the class (50) gives rise to ghost and asymptotically free theories.
 Below we shall analyze the dynamics of these theories in the metric-affine approach.
@@
 Indeed, despite the complexity of the metric field equations, the connection field equations obtained above, have a rather simple structure.
-As a matter of fact, since the form of (45) is identical to that of (14) with only the exact form of $\phi$ altered, the exact and unique solution to (45) is still
+As a matter of fact, since the form of (47) is identical to that of (14) with only the exact form of $\phi$ altered, the solution to (47) is still
 $N_{\alpha\mu\nu}=\frac{2}{(n-2)\phi}g_{\nu[\alpha}\partial_{\mu]}\phi+\frac{1}{2}g_{\alpha\mu}q_{\nu}$ (49)
@@
 IV.2 Non-local Infrared $\mathcal{L}=R+RF(\Box^{-1})R$ Theories
-The previous results can be straightforwardly applied to non-local theories involving the inverse box operator Deser:2007jk $\Box^{-1}$ defined through $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$.
+The previous results can be straightforwardly applied to non-local theories involving the inverse box operator Deser and Woodard (2007) $\Box^{-1}$ defined through $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$.
 As a prototype let us consider the class of non-local theories given by [^8] For the study of such theories in the metric formalism see e.g.
-Capozziello:2020xem ; Capozziello:2021bki ; Capozziello:2023mju .
+Capozziello et al.
+(2020); Capozziello and Capriolo (2021); Capozziello et al.
+(2023).
 $S[g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big(R+RF(\Box^{-1})R\Big)$ (50)
@@
 Note that this form of series for $F$, namely containing powers of the inverse box operator are relevant for the IR regime.
-In order to proceed with the field equations, we only need to vary the defining relation of $\Box^{-1}$ and subsequently multiply from the left by the same operator, to trivially get
+In order to proceed with the field equations, we only need to vary the defining relation of $\Box^{-1}$ and subsequently multiply from the left by the same operator, to trivially get [^9] We assume that a Green’s function and boundary conditions are fixed and no homogeneous modes of $\Box$ appear.
 $\delta\Box^{-1}=-\Box^{-1}(\delta\Box)\Box^{-1}$ (52)
@@
 Therefore, in a MAG scheme before extending to (11) it would be quite natural to consider non-local actions consisting of scalars formed by sandwiching the box operator with torsion and non-metricity before moving to quadratic curvature.
-Given that there are 11 parity even quadratic invariants, the analogue of the theory (11) to MAG is
+Given that there are 11 parity even quadratic invariants (see Pagani and Percacci (2015) and also Iosifidis and Koivisto (2019) for the proof that this is the number of independent invariants), we get the following scalars
+Pure Non-Metricity Scalars
+$A_{1}=Q_{\alpha\mu\nu}Q^{\alpha\mu\nu}$ (56)
+$A_{2}=Q_{\alpha\mu\nu}Q^{\mu\nu\alpha}$ (57)
+$A_{3}=Q_{\mu}Q^{\mu}$ (58)
+$A_{4}=q_{\mu}q^{\mu}$ (59)
+$A_{5}=Q_{\mu}q^{\mu}$ (60)
+Pure Torsion Scalars
+$B_{1}=S_{\alpha\mu\nu}S^{\alpha\mu\nu}$ (61)
+$B_{2}=S_{\alpha\mu\nu}S^{\mu\nu\alpha}$ (62)
+$B_{3}=S_{\mu}S^{\mu}$ (63)
+Mixed
+$C_{1}=Q_{\alpha\mu\nu}S^{\alpha\mu\nu}$ (64)
+$C_{2}=Q_{\mu}S^{\mu}$ (65)
+$C_{3}=q_{\mu}S^{\mu}\,$, (66)
+and then the analogue of the theory (11) to MAG is
 $S_{[}g,\Gamma]=\frac{1}{2\kappa}\int d^{n}x\sqrt{-g}\Big[R+b_{1}S_{\alpha\mu\nu}S^{\alpha\mu\nu}+b_{2}S_{\alpha\mu\nu}S^{\mu\nu\alpha}+b_{3}S_{\mu}S^{\mu}+c_{1}Q_{\alpha\mu\nu}S^{\alpha\mu\nu}+c_{2}Q_{\mu}S^{\mu}+c_{3}q_{\mu}S^{\mu}$
@@
 These exhaust the number of parity even invariants up to dimension $L^{-2}]$.
-When only the first two lines of the above action are considered then the vacuum theory reduces to General Relativity Iosifidis:2021tvx .
+When only the first two lines of the above action are considered then the vacuum theory reduces to General Relativity Iosifidis (2022a).
 The third and fourth lines are the new non-local terms we consider.
@@
 where $N$ is the distortion.
-This shows that torsion and non-metricity propagate in this case and the theory has more degrees of freedom than General Relativity.
+This shows that torsion and non-metricity have dynamics in this case and the theory has more degrees of freedom than General Relativity.
 It goes beyond the scope of the current study to investigate the healthy sub-sectors of this class of theories.
@@
 This is of course tightly related with the fact that two connections are related through a projective transformation and have their symmetric Ricci tensor equal.
-Furthermore the connections, in this equivalence class, share also the same autoparallel curves up to a reparameterization of the affine parameter Iosifidis:2018zjj .
+Furthermore the connections, in this equivalence class, share also the same autoparallel curves up to a reparameterization of the affine parameter Iosifidis et al.
+(2019).
 Finally, varying with respect to the connection and using the form of (72) we find the metric field equations
@@
 Of course, it goes without saying that in general, these excitations extra degrees of freedom will lead to pathologies.
-However, it should be noted that the additional terms we consider here when added to the quadratic MAG theory could extend the spectrum of healthy theories just like the inclusions of cubic terms in the action, see Bahamonde:2024efl .
+However, it should be noted that the additional terms we consider here when added to the quadratic MAG theory could extend the spectrum of healthy theories just like the inclusions of cubic terms in the action, see Bahamonde and Gigante Valcarcel (2025).
 VI Applications to Cosmology
@@
 For completeness, we will also report the corresponding cosmological equations for the novel theory (69).
-Let us recall that, in this case, since only the vector part of torsion is excited, in the FLRW spacetime the full torsion is fully described by a single scalar field and the torsion vector takes the form tsamparlis1979cosmological
+Let us recall that, in this case, since only the vector part of torsion is excited, in the FLRW spacetime the full torsion is fully described by a single scalar field and the torsion vector takes the form Tsamparlis (1979)
 $S_{\mu}=3\Phi u_{\mu}$ (83)
@@
 Acknowledgements
-The authors acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), Sezioni di Napoli e di Torino, Iniziative Specifiche QGSKY and MOONLIGHT2.
+We thank Friedrich H.
+Hehl for useful discussions and comments.
+The authors acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Napoli, Iniziative Specifiche QGSKY and MOONLIGHT2, and the Istituto Nazionale di Alta Matematica (INdAM), Gruppo Nazionale di Fisica Matematica (GNFM).
 This paper is based upon work from COST Action CA21136 Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by COST (European Cooperation in Science and Technology).
@@
 The variations for the first two lines in (67) are well known, see e.g.
-Iosifidis:2021tvx , so we will concentrate on the derivation of the new part.
+Iosifidis (2022a), so we will concentrate on the derivation of the new part.
 Set
@@
 The $g$-variation first two lines that contain no box operator is well known, see e.g.
-Iosifidis:2021bad .
+Iosifidis (2022b).
 The lines containing the box operator are trickier so let us elaborate on them in detail.